OpenStudy (anonymous):

The coffee shop found another delivery company that sells orders at increments of 500 cups, Coffee Delivery B. They charge $3.50 each week to be on their delivery route and charge 3.9 cents per disposable cup.

OpenStudy (anonymous):

How do you make a function using the information about the second delivery company? @jim_thompson5910

OpenStudy (anonymous):

Is it f(x) = 0.039x + 3.50?

jimthompson5910 (jim_thompson5910):

close, you forgot to factor in the fact that they sell "orders at increments of 500 cups"

OpenStudy (anonymous):

So how do i do that @jim_thompson5910

OpenStudy (doc.brown):

You have 2 rates: 500 cups = 1 increment $0.039 = 1 cup And another rate added on: 1 week = $3.50

OpenStudy (anonymous):

So, f(x) = 0.039x + 3.50c? I would have to graph this too.

OpenStudy (doc.brown):

I would do $ per increment plus the weekly rate.

OpenStudy (anonymous):

Can you show me how the function will look like?

OpenStudy (doc.brown):

I'll show you the conversion first.\[\frac{500 cups}{increment}\times\frac{$0.039}{cup}+\frac{$3.50}{week}\\\frac{500\cancel{cups}}{increment}\times\frac{$0.039}{\cancel{cup}}+\frac{$3.50}{week}\\\frac{$19.50}{increment}+\frac{$3.50}{week}\]

OpenStudy (anonymous):

Oh I see how you perceive it. Well my function needs to consist of 0.039x or else the teacher might count it as wrong.

OpenStudy (doc.brown):

If the company has 1 customer that buys 0 cups, they will make $3.50 profit per week. The function is dependent on 2 variables, number of customers and number of increments sold. I would plot it as a function of time, profit over weeks.

OpenStudy (anonymous):

ok...so how would i graph it?

OpenStudy (doc.brown):



OpenStudy (anonymous):

Ok now I understand. Thank you so much!

OpenStudy (doc.brown):

I am pretty sure your teacher may have something different in mind. Maybe fish for another opinion to consider over mine?

OpenStudy (anonymous):

Yeah but I'll try to combine your way and my way, and I'll see how that goes.